its a word problem :
michelle is 5 years younger than cindy . janet is ten years less than twice of cindy age . kenny is 10 years old . the ratio of janets age equal s the ratio of michelles age to kenny's age . how old are cindy , janet, and michelle ?

Answer by palanisamy(496)

The age of Kenny = 10
Let the age of Cindy be x
Janet is ten years less than twice of cindy age
The the age of Janet = 2x-10
Michelle is 5 years younger than cindy
The the age of Michelle = x-5
Given, the ratio of janets age equal s the ratio of michelles age to kenny's age
2x-10 = (x-5)/10
20x-100 = x-5
20x-x = 100-5
19x = 95
x = 95/19
x = 5
Therefore the age of Cindy = 5
the age of Janet = 2*5-10 = 0
the age of Michelle= x-5 = 5-5 = 0